Session: Feb 1999 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K.Borkowski at Tue Feb 16 09:37:15 MET 1999) Success
No problems reported.

Onsala (A. Polatidis at Tue Feb 16 12:11:31 MET 1999) Minor failures
No data after 1924 UT. Antenna had to be stowed due to high winds, which reached up to 17m/s.

Noto (C.Trigilio at Tue Feb 16 18:25:50 MET 1999) Success

Medicina (Bondi at Wed Feb 17 09:18:54 MET 1999) Success
No known problem

Westerbork (Tony Foley at Wed Feb 17 13:54:06 MET 1999) Success
No known problems

Jodrell (Alastair Gunn at Fri Feb 19 14:12:52 MET 1999) Minor failures
Telescope parked at 00:39 UT due to high winds. Did not resume. No other known problems.

Shanghai (Huang Xinyong at Mon Feb 22 10:57:42 MET 1999) Success
It is cloudy.

Effelsberg (dgraham at Mon Feb 22 15:24:37 MET 1999) Minor failures
stopped for snow 1640-1730. Tsys near end variable due to thick cloud.

Calibration (Jean\-Francois Desmurs at Wed Mar 3 11:28:14 MET 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the feb99 directory under the name: v085c.antab. Read the comments in the file to get further information.